OpenStudy (anonymous):

The volume of a cone is 300 cubic kilometers. If the radius of the base is 7 kilometers, what is the height of the cone? Use for 3.14 . Show your work and round to the nearest tenth.

OpenStudy (anonymous):

Do you know the formula of the cone volume?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

can you write it?

OpenStudy (anonymous):

because you have the volume and pi = 3.14 and the radius, so you can figure out the height

OpenStudy (anonymous):

v=(1/3) pi r^2 h

OpenStudy (anonymous):

ok 300 = (1/3) * (3.14) * (7*7) h so h = (300*3) / (3.14*7*7) h = 900/153.86 h = 5.85 Kilometer

OpenStudy (anonymous):

get it @CMangini2 ??

OpenStudy (anonymous):

Yeah Thankn you

OpenStudy (anonymous):

No problem anytime!